OpenStudy (anonymous):

Factor the Expression 8 - m^6

OpenStudy (anikhalder):

Key: ^ denoting 'to the power of' 8=2^3 m^6=(m^2)^3 let (m^2) =p so, (m^6) = p^3 So, 8-m^6 =2^3 - p^3 Use the formula: (a^3 - b^3)= (a-b)(a^2 +ab +b^2) So we can further factorize the initial expression